OpenStudy (anonymous):

find the equation of thw line through the point (-2,-2) that is perpendicular to the line y = 2x - 1

OpenStudy (anonymous):

y=x/2-1

OpenStudy (anonymous):

since it's prependicular to the line y=2x-1 therfore m1*m2=-1 where m is the slope . therfore m2 = -1/2 the equation may be writen as (y-y1)/(x-x1)=m in other words (y+2)/(x+2)=-1/2 y=-1/2x-1

OpenStudy (anonymous):

sorry it will be (-x/2) -3